OpenStudy (anonymous):

What is the inverse of the function y=-x-1

jimthompson5910 (jim_thompson5910):

Swap x and y to go from y = -x-1 to x = -y-1

jimthompson5910 (jim_thompson5910):

Then solve for y

OpenStudy (anonymous):

I did and I got the same as before I did the inverse... is that correct?

jimthompson5910 (jim_thompson5910):

that is correct, the reason why is because y = -x-1 is perpendicular to y = x

jimthompson5910 (jim_thompson5910):

if you reflect y = -x-1 over the line y = x, you'll see that the line y = -x-1 will reflect over itself (ie it's symmetrical about the line y = x)

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

np